Mr. Fairbanks invested half his money in land, a tenth in stock, and a twentieth in bonds. He put the remaining $35,000 in a savings account. What is the total amount of money that Mr. Fairbanks saved or invested?
Answer by rfer(16322)

1.0-(.5+.1+0.083333)=0.
35000/0.316667=$110,526
------------------
1/2+1/10+1/12=
60/120+12/120+10/120=82/120=41/60
60/60-41/60=19/60
35000*60/19=2100000/19=$110,526